OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular to the line y = three-fifths x minus 1 and passes through the point (-9, 4).

OpenStudy (ivettef365):

so you have y = 3/5x -1

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

to be perpendicular the slope needs to be opposite reciprocal. what is the slope on that equation

OpenStudy (anonymous):

5/3 ?

OpenStudy (ivettef365):

the slope on that equation is actually 3/5 y = mx + b where m = slope

OpenStudy (ivettef365):

now what is opposite reciprocal of 3/5

OpenStudy (anonymous):

5/3

OpenStudy (ivettef365):

the reciprocal is 5/3 but opposite reciprocal = -5/3

OpenStudy (anonymous):

oh

OpenStudy (ivettef365):

now you need to find the equation of the line that goes through points (-9,4) you use y = mx + b you have m = -5/3 you have x = -9 you have y = 4 so the only thing you need to find is b can you set it up to find b

OpenStudy (anonymous):

I honestly have no idea

OpenStudy (ivettef365):

ok you have y = mx + b, to set up just substitute letters for the numbers it would look like this 4 = -5/3(-9) + b

OpenStudy (ivettef365):

now to solve for b as follows: 4 = (-5) (-3) + b 4 = 15 + b 4 - 15 = b -11 = b so b = -11

OpenStudy (ivettef365):

now the equation of the line perpendicular that passes through those points is y = -5/3x -11

OpenStudy (anonymous):

thanks ivettef365

OpenStudy (ivettef365):

yw :)